samantha charges $27.28 for 14 tacos and 6 buritos and $18.52 for 8 tacos and 6 burritos determine the cost of one burrito and one taco: samantha charges $27.28 for 14 tacos and 6 buritos and $18.52 for 8 tacos and 6 burritos determine the cost of one burrito and one taco
Answer by nerdybill(1040)

samantha charges $27.28 for 14 tacos and 6 buritos and $18.52 for 8 tacos and 6 burritos determine the cost of one burrito and one taco
.
Assign variables to things you don't know:
Let t = cost of a taco
and b = cost of a burrito
.
Since you have two unknowns, you'll need to find two equations.
.
From: "samantha charges $27.28 for 14 tacos and 6 burritos" we get equation 1:
14t + 6b = 27.28
.
From: "$18.52 for 8 tacos and 6 burritos" we get equation 2:
8t + 6b = 18.52
.
The two equations are:
14t + 6b = 27.28
8t + 6b = 18.52
.
If we were to multiply both sides of equation by a -1 we'll get:
14t + 6b = 27.28
-8t - 6b =-18.52
.
We can now ADD the two equations to eliminate b's:
14t + 6b = 27.28
-8t - 6b =-18.52
-----------------
6t = 8.76
Dividing both sides by 6:
t = $1.46 (cost of a taco)
.
Plug the above into equation 2 and solve for b:
8t + 6b = 18.52
8(1.46) + 6b = 18.52
11.68 + 6b = 18.52
6b = 6.84
b = $1.14 (cost of a burrito)